AZ Q Science night

0:10 Josh brough up wavefront sensor in the afternoon. He noted chevron pattern, followed by OCAM not live, so decided to just use the hammer to: power off, restart the software, power on again.

Talked to Josh about wfs_init to be added to the OSA procedure rather than W_stop_all/Start_all. He is putting that in.

2:11. OCAM went not live during AO preset for OSCO and (2:10:26)
WfsArbitrator_129758|ERR| 171|2021-04-05 02:10:26.955081| MAIN > exception ::AODefs::AOException
WfsArbitrator_129758|ERR| 172|2021-04-05 02:10:26.955168| MAIN > {
WfsArbitrator_129758|ERR| 173|2021-04-05 02:10:26.955200| MAIN > code = -1
WfsArbitrator_129758|ERR| 174|2021-04-05 02:10:26.955234| MAIN > what = [AOException] /home/AOeng/aoroot/calib/wfs/current/ocam2/LUTs/current/pixelGating.txt (code -1010) File open/read/write error File BCUlbt.cpp line 485

Called greg, but silly me didn't realize it was just a reversion of the ACE table. Thanks Greg!!

There was something wrong with the new bin1 file, so he reverted to the previous one. Bin 1: 20210210_205151

Resent preset right after he fixed the ACE table but....
WfsArbitrator_129758|ERR| 225|2021-04-05 02:24:07.331277| MAIN > Attaching to shared buffer mirrorctrl.L:TECHVIEWER_65536 (-3509) BUF_NOT_AVAIL_ERROR
WfsArbitrator_129758|DEB| 226|2021-04-05 02:24:07.331763| MAIN > arbFunc updateModulationInternal (, 'FLAOWFS', 'LUCIFER', 200)
WfsArbitrator_129758|INF| 227|2021-04-05 02:24:07.332136| MAIN > Tip-tilt frequency is not synced, skipping update
WfsArbitrator_129758|DEB| 228|2021-04-05 02:24:07.332188| MAIN > arbFunc updateModulationInternal done.

Cycled the WFS ARB. I should have done that after the ACE table update. CCD47 not live so cyced BCU and rebinned. Everything happy now!

2:32 OSCO preset sent successfully in AUTO. HD71514. WFS Estimated mag 9.7. Optical Gain 0.327

Not executing the NCPA on because we are running pretty late. Time for science.

2:43 Off to an ESM science script. AOref catalog mag 14.25. Seeing 0.8". Some cirrus. Adjusted frequency down a touch since we were losing photons to clouds and the average pupil intensity was very low. We can adjust to a lower binning if this doesn't hold, but it does look good.

3:08. Camera lens is railed, and we are not meeting the PI's science requirements: 0.77pixel off.

PI requires <0.4" on science images

Given overheads of resetting up for a different PA or binning to get the camera lens range, and that conditions were not condusive with their science at this time we are moving on to another high priority science target (seeing limited LUCI).

6:07 I neglected to put in the Adsec 26 degree elevation limit at the start of the night. Implementing now.

8:36 preset to new ESM target from SL target at 26deg. I noticed that the set shape occurred while we were very low

ESM targ with very faint AO ref 16.78 catalog mag, 17.3 estimated magnitude on the WFS. Seeing 0.7" on the guiders (luci2 running in parallel seeing limited). Bin4, frequency 200, lots of range on the camera lens with a PA of 178 and a rotator angle of 70deg and rising.

TN# 20210405_084202

10:11 Preset to ESM targ (LUCI2 running in SL). AO ref 15.35 catalog mag. RIP on optimize gain, rotator 422

Recover fail at 47 degrees, flat set with 447 modes

Had Josh unwrap the rotator as well to give us rerotator range, now at 63. pupil difference ok but Y position is nearly ~98-100
TN#20210405_102435

Catalog magnitude 15.24, estimated WFS magnitude 16.0, bin4, 400 modes. 5.8 photons/subapp, ~30 average pupil intensity.

10:44 Cameralens is backing away from saturation, rotator at 69. TN#202105405_104550

12:13 All operations cdone for tonight. WFS warmed an powered off

-- JenniferPower - 05 Apr 2021
Topic revision: r6 - 05 Apr 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback